OpenStudy (anonymous):

Polynomial powered by 2

OpenStudy (anonymous):

A polynomial f(x) has degree 8 and \[\huge f(i)=2^i\] for\[ i=0,1,2,3,4,5,6,7,8\]. Find \\[\huge \color{cyan}{f(9)}\].

OpenStudy (anonymous):

@campbell_st @phi @TuringTest @shubhamsrg @sauravshakya @satellite73

OpenStudy (anonymous):

Isn't it solved using the lagrange interpolating polynomial? I'll study it in a couple of days, and I'll be able to help you without googling it xD

OpenStudy (anonymous):

f(x)=(x-1)(x-2)(x-3)(x-4)(x-5)(x-6)(x-7)(x-8)/((-1)(-2)(-3)(-4)(-5)(-6)(-7)(-8))+2(x)(x-2)(x-3)(x-4)(x-5)(x-6)(x-7)(x-8)/((1)(-1)(-2)(-3)(-4)(-5)(-6)(-7))+4(x)(x-1)(x-3)(x-4)(x-5)(x-6)(x-7)(x-8)/((2)(1)(-1)(-2)(-3)(-4)(-5)(-6))+8(x)(x-1)(x-2)(x-4)(x-5)(x-6)(x-7)(x-8)/((3)(2)(1)(-1)(-2)(-3)(-4)(-5))+16(x)(x-1)(x-2)(x-3)(x-5)(x-6)(x-7)(x-8)/((4)(3)(2)(1)(-1)(-2)(-3)(-4))+32(x)(x-1)(x-2)(x-3)(x-4)(x-6)(x-7)(x-8)/((5)(4)(3)(2)(1)(-1)(-2)(-3))+64(x)(x-1)(x-2)(x-3)(x-4)(x-5)(x-7)(x-8)/((6)(5)(4)(3)(2)(1)(-1)(-2))+128(x)(x-1)(x-2)(x-3)(x-4)(x-5)(x-6)(x-8)/((7)(6)(5)(4)(3)(2)(1)(-1))+256(x)(x-1)(x-2)(x-3)(x-4)(x-5)(x-6)(x-7)/((8)(7)(6)(5)(4)(3)(2)(1))

OpenStudy (anonymous):

man this is hectic you know its for hish school students i bet they have to know lagrange

OpenStudy (anonymous):

@shubhamsrg @UnkleRhaukus

OpenStudy (shubhamsrg):

Stuck on as this also.

OpenStudy (anonymous):

it looks like the equation has to behave like an exponential function 2^i hence it has to be divisible by 2 @ SqueeSpleen this numbers i=0,1,2...8 are not roots

OpenStudy (shubhamsrg):

I got the answer. I have a logic also but it is not perfect, still, I am able to get the correct answer. Answer is 511.

OpenStudy (anonymous):

i only se that the constant is 1 and the closest function that is like 2^x i got is \[f(x)=(\frac{ x }{ 10 }+1)^8\]

OpenStudy (shubhamsrg):

Here is what I did I started with lower degrees, For each, I took for granted that f(i) = 2^i , where 0<= i <=n (n is the degree) after finding f(x) with degree n, I found out f(n+1) Note that this is all symmetrical to what is asked in the actual ques. for degree 1, let polynomial be ax+b so a(0) + b =1 and a(1) + b = 2 So our polynomial is x+1 , f(2) = 3 for degree 2, let polynomial be ax^2 + bx + c so a(0) + b(0) + c =1 a(1) + b(1) + c = 2 a(4) + b(2) + c = 4 So our polynomial is (1/2)(x^2 +x +2) f(3) = 7 Same way when you solve for degree 3, f(4) = 15 There is a pattern 3 , 7, 15 , next term should then be 31 owing to the pattern. then 63 , 127,255, 511 511 comes when degree is 9. And to my surprise, it is the right answer.

OpenStudy (shubhamsrg):

I still don't have any idea why this pattern is followed or, what is the correct procedure for the question. @mukushla

OpenStudy (anonymous):

this is a good method and the fact that it anwsers the question is enough the pattern is very true

OpenStudy (anonymous):

The numbers 0,..,8 aren't roots in my equations, it has 9 terms, and when you evaluate the equation in i, the terms 0..,i-1 and i+1,...,8 become zero, but the term i becomes f(i) I didn't know this was for high school students :P

OpenStudy (anonymous):

And I forget to evaluate the polynomial in f(9) xD

OpenStudy (anonymous):

Yes, the answer is 511.

OpenStudy (anonymous):

The polynomial simplified is: (40320 + 25584 x + 15980 x^2 - 4340 x^3 + 3969 x^4 - 1064 x^5 + 210 x^6 - 20 x^7 + x^8)/40320 I know the question wasn't what's the polynomial, but what's the value of f[9], but I forget the question because I read a little about it a few days ago xD

OpenStudy (anonymous):

@mukushla